OpenStudy (anonymous):

The graph of which quadratic equation is shown below? a u-shaped graph on a coordinate plane that is opening down and has a vertex of (3, 4) y = x2 + 6x − 5 y = x2 − 6x − 5 y = −x2 − 6x − 5 y = −x2 + 6x − 5

OpenStudy (anonymous):

help me please

OpenStudy (tkhunny):

Opening down is a good clue. Keep all the -x2 and discard all the +x2. What do you know about Completing the Square?

OpenStudy (anonymous):

nothing

OpenStudy (tkhunny):

Not sure how to help you, then. It is an important skill that you should have been taught before facing this problem. Would like to see it, once? Then you can go complain to your teacher for not giving you the tools you need to solve the problems you have been given.

OpenStudy (anonymous):

okay

OpenStudy (tkhunny):

Actually, we can solve this one without that - sort of reverse completing the square. Downward pointing parabolas with vertex at (h,k) look something like this: \(y - k = -(x-h)^{2}\) Substituting the know vertex, we get \(y - 4 = -(x-3)^{2}\) Expanding \(y - 4 = -(x^{2} - 6x + 9)\) Removing parentheses \(y - 4 = -x^{2} + 6x - 9\) Solving for y \(y = -x^{2} + 6x - 5\) Did we find it?

OpenStudy (anonymous):

nope

OpenStudy (tkhunny):

Where are you looking? Isn't that the last one? :-)

OpenStudy (anonymous):

wait i didnt see it

OpenStudy (anonymous):

yep i saw it

OpenStudy (anonymous):

thank you